OpenStudy (buckybarnes):

Find the distance between points M(6,16) and Z(-1,14) to the nearest tenth

OpenStudy (buckybarnes):

I got 5.4

OpenStudy (buckybarnes):

@3mar

OpenStudy (buckybarnes):

@Wholock

OpenStudy (3mar):

\[\sqrt{(x_2^{2}-x_1^{2})+(y_2^{2}-y_1^{2})}\] I think you know that expression. don't you?

OpenStudy (buckybarnes):

Yes

OpenStudy (3mar):

Simply substitute: You got: \[\sqrt{(6+1)^2+(16-14)^2}\] \[=\sqrt{53}=7.28\] please, check my result and yours to be sure.

OpenStudy (buckybarnes):

Would it be 7.3 since I have to round it to the nearest tenth?

OpenStudy (buckybarnes):

@3mar

OpenStudy (3mar):

Yes, but are you convinced?

OpenStudy (buckybarnes):

Convinced?

OpenStudy (3mar):

Thanks for the medal, by the way.

OpenStudy (3mar):

"completely certain about something"

OpenStudy (skullpatrol):

ie informally prove it to yourself, for example, by drawing a scale diagram and measuring it with ruler?

OpenStudy (skullpatrol):

that^ will help you feel certain that your answer is correct :)